# Quantum Adaptive Self-Attention for Quantum Transformer Models

### v2

@@ Abstract
Transformer models have revolutionized sequential learning across various domains, yet their self-attention mechanism incurs quadratic computational cost, posing limitations for real-time and resource-constrained tasks. To address this, we propose Quantum Adaptive Self-Attention (QASA), a novel hybrid architecture that enhances classical Transformer models with a quantum attention mechanism. QASA replaces dot-product attention with a parameterized quantum circuit (PQC) that adaptively captures inter-token relationships in the quantum Hilbert space. Additionally, a residual quantum projection module is introduced before the feedforward network to further refine temporal features. Our design retains classical efficiency in earlier layers while injecting quantum expressiveness in the final encoder block, ensuring compatibility with current NISQ hardware. Experiments on synthetic time-series tasks demonstrate that QASA achieves faster convergence and superior generalization compared to both standard Transformers and reduced classical variants. Preliminary complexity analysis suggests potential quantum advantages in gradient computation, opening new avenues for efficient quantum deep learning models.
@@ 1 Introduction
Transformer architectures [VSP+17] have emerged as the backbone of modern deep learning, powering state-of-the-art models in natural language processing, vision, and sequential prediction. Their core strength lies in the self-attention mechanism, which models long-range dependencies dynamically. However, this comes at a significant computational cost—quadratic in sequence length—hindering real-time applications and large-scale deployment.
Meanwhile, quantum computing has introduced new paradigms in learning representations through entanglement and high-dimensional Hilbert spaces. Motivated by this, we explore the fusion of quantum computing and Transformers, focusing on enhancing attention mechanisms with quantum adaptability.
In this work, we introduce Quantum Adaptive Self-Attention (QASA)—a hybrid classical-quantum attention module designed to replace traditional dot-product operations with a learnable parameterized quantum circuit. Unlike purely quantum or classical models, QASA offers a practical balance: classical encoder layers ensure stable training, while a quantum-enhanced final layer boosts expressiveness and captures non-classical correlations.
1.1 Our contributions
  • Quantum Adaptive Self-Attention Transformer Block (QASA): We propose a quantum residual attention mechanism integrated into a Transformer architecture.
  • Quantum Feature Enhancement Layer: We introduce a quantum projection layer that refines sequence features prior to final prediction.
  • We empirically demonstrate the effectiveness of QASA in time-series forecasting, supported by theoretical insights on quantum complexity benefits.
Figure 1: QASA model Architecture.
@@ 2.1 Deep Learning on Time-Series Data
Deep learning has significantly advanced the modeling and understanding of time-series data across various domains. Recurrent Neural Networks (RNNs) [SP97], particularly Long Short-Term Memory (LSTM) networks [HS97] and Gated Recurrent Units (GRUs) [CGCB14], have been widely adopted due to their ability to capture long-term temporal dependencies. However, their sequential nature limits parallelization and increases training time. To address this, temporal convolutional networks (TCNs) [LFV+17] and Transformer-based architectures [VSP+17] have demonstrated superior performance by enabling parallel processing and better long-range dependency modeling.
Recent work in biomedical domains, such as electroencephalography (EEG), has utilized deep learning models to capture non-linear temporal dynamics in brain activity. Convolutional neural network-Transformer hybrids and attention-based models have shown success in tasks like motor imagery classification [CW24a], seizure detection [CCT25b], response prediction [LCC+23] and visual stimuli estimation [CW24b, Che24]. In industrial applications, particularly supply chain demand forecasting, Transformer variants such as the Temporal Fusion Transformer (TFT) [LALP21] and Informer [ZZP+21] have been applied to multi-horizon prediction tasks with irregular time intervals and exogenous variables. Similar time series learning techniques have been applied successfully in Molecule dynamics and even open quantum systems [TFX+22, TKT20].
Despite these advancements, challenges remain in modeling noisy, sparse, or irregularly-sampled sequences—especially in domains like healthcare and logistics—motivating [CC25] the development of hybrid models that integrate domain-specific priors with general-purpose deep learning architectures.
@@ 2.2 Quantum Deep Learning on Time-Series Data
While classical deep learning has achieved remarkable success in time-series analysis, it often requires extensive computational resources and large-scale data to generalize effectively. With the advent of quantum computing, Quantum Deep Learning (QDL) [BWP+17] has emerged as a promising paradigm for modeling complex temporal patterns, offering theoretical advantages in expressivity and optimization through quantum entanglement and superposition.
In the context of time-series data, several quantum-inspired architectures have been proposed. Variational Quantum Circuits (VQCs) [GJ24] and Quantum Recurrent Neural Networks (QRNNs) [TMN+21] have been introduced to capture temporal dependencies using parameterized quantum gates and recurrent structures. A particularly notable development is the Quantum Long Short-Term Memory (QLSTM) network [CYF22], which integrates quantum circuits into the gating mechanisms of classical LSTMs. By replacing linear transformations with quantum operations, QLSTM models can potentially enhance memory capacity and temporal sensitivity, even in low-data regimes.
Applications of these models span multiple domains. In EEG signal analysis, hybrid QNN and quantum convolutional models have demonstrated improved efficiency in capturing spatiotemporal brainwave dynamics for tasks like motor imagery classification [CCTW24, CCT25a] and cognitive state detection [CTH24]. In finance and supply chain forecasting, quantum-enhanced LSTM architectures have been investigated for demand prediction and volatility modeling, leveraging quantum parallelism for more efficient scenario sampling and non-linear dynamics modeling [PS22].
Despite these advancements, research on Quantum Transformers for time-series data remains limited. Most existing studies focus on Quantum Visual Transformers (QViT) [CKM+22, KMCC24] in image-based tasks, leaving a gap in understanding how quantum self-attention mechanisms can be leveraged for sequential data. Addressing this gap is the key focus of this work. By exploring quantum self-attention architectures tailored to temporal patterns, we aim to provide novel insights and practical tools for efficient and expressive time-series modeling in quantum machine learning.
@@ 3 Methodology
Figure 2: The variational quantum circuit (quantum neural network) used in this work.
Table 1: Quantum Adaptive Self-Attention (QASA) Transformer Architecture. Each row lists a layer with its operation and input/output tensor shapes. $L$ denotes the input sequence length (e.g., $L{=}50$), $d$ is the hidden feature dimension (e.g., $d{=}256$), and $n$ is the number of qubits in the quantum circuit (e.g., $n{=}8$). ‘Attn’ is an abbreviation for ‘attention’.
@@
Transformer Layer ×$\times(N{-}1)$ Multihead Attn + FFN $(L,d)$ $(L,d)$
@@ 3.3 Transformer Encoder Layers
The encoder consists of $N$ layers, where the first $N-1$ layers are standard Transformer encoder layers defined as:
@@
[…]
@@ 3.4 Quantum Encoder Layer
where $\text{where }head_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}),$ (15)
@@
The vector $h_{q}$ is then passed as input to a parameterized quantum circuit (PQC), which is defined over $n+1$ qubits and $L_{q}$ layers of unitary operations. The PQC applies data re-uploading using $RX$ and $RZ$ gates, and introduces entanglement using $CNOT$ and $RY$ gates. The final output is computed as the expectation values of Pauli-Z operators on the first $n$ qubits:
@@
The resulting vector $h_{q}$ is fed into a parameterized quantum circuit $\text{QC}(\cdot)$ with $L_{q}$ layers and $n+1$ qubits, implemented using a data reuploading strategy. Each layer of the PQC applies a combination of single-qubit rotations ($RX$, $RZ$, $RY$) and entangling gates (e.g., $CNOT$), allowing the circuit to perform complex non-classical transformations:
@@
In our design, we adopt a hybrid encoder composed of $(N{-}1)$ standard Transformer encoder layers followed by a single quantum-enhanced encoder layer. This layered configuration offers a practical and effective trade-off between scalability and expressiveness. Specifically, the classical Transformer blocks serve as powerful hierarchical feature extractors with stable training dynamics, while the quantum encoder layer provides a complementary inductive bias via entangled quantum operations and high-dimensional projection.
@@
(d) $QASA_{classical}$ Epoch 1
(e) $QASA_{classical}$ Epoch 15
(f) $QASA_{classical}$ Epoch 30
(g) $QASA$ Epoch 1
(h) $QASA$ Epoch 15
(i) $QASA$ Epoch 30
@@
    Data Encoding via Single-Qubit Rotations: The input features are encoded into quantum states using $RX$ and $RZ$ gates per qubit:
    $\forall i\in\{0,\dots,n-1\},\quad\text{Apply }RX(x_{i}),\ RZ(x_{i})\text{ on wire }i.$ (25)
@@
    Parameterized Rotations with Data Reuploading: Across $L_{q}$ layers, the circuit includes learnable $RY$ and $RZ$ rotations:
    $RY(\theta_{l,i})\ RZ(\theta_{l,i})\quad\forall i,\forall l\in\{1,\dots,L_{q}\},$ (26)
@@
    $\text{CNOT}(i\rightarrow(i+1)\mod n),\quad\text{for }i=0\dots n{-}1.$ (27)
@@
    $\text{CNOT}(n{-}1\rightarrow n),\quad RY(\theta_{l,n})\text{ on wire }n.$ (28)
This gate configuration balances expressiveness and hardware feasibility. The use of simple universal gates ($RX,RZ,RY$, CNOT) ensures compatibility with most current quantum hardware, while circular entanglement introduces full connectivity with only $n$ CNOTs per layer, avoiding unnecessary depth. The auxiliary $(n+1)$-th qubit also serves as a tunable global channel, useful for capturing long-range dependencies in the encoded features.

### v3

@@ Abstract
[…]
keywords:
quantum machine learning, quantum transformer, self-attention, time-series forecasting, NISQ, hybrid quantum-classical
articletype: Paper
@@ 1 Introduction
Transformer architectures [39] have become the backbone of modern deep learning, powering state-of-the-art models across natural language processing, computer vision, and sequential prediction. Their core strength lies in the self-attention mechanism, which dynamically models long-range dependencies. However, this comes at a significant computational cost—quadratic in sequence length—motivating the search for alternative attention paradigms.
Quantum computing offers a compelling candidate: parameterized quantum circuits (PQCs) operate in exponentially large Hilbert spaces and can represent correlations that are classically intractable to express compactly [22]. Recent work has explored quantum-enhanced sequence models, including quantum LSTMs [15] and quantum Transformer variants [5], which replace classical components wholesale with variational quantum circuits. However, a fundamental question remains unanswered: how much quantum integration is actually beneficial, and where should it be placed?
In this work, we propose Quantum Adaptive Self-Attention (QASA), a hybrid architecture designed around a principle of architectural parsimony—the idea that maximal quantum benefit comes not from maximizing quantum resources, but from placing minimal quantum computation at the optimal position. QASA retains N1N{-}1 classical Transformer encoder layers and replaces only the value projection in the final layer with a PQC, using just 36 trainable quantum parameters. This stands in contrast to approaches such as QLSTM (128 quantum parameters) and QnnFormer (90 quantum parameters), which distribute quantum computation across the entire model.
Our systematic evaluation across nine synthetic benchmarks, one real-world dataset, two quantum baselines (QLSTM and QnnFormer) alongside a classical Transformer, and a comprehensive ablation study yields the following key contributions:
1.1 Contributions
[…]
  • Task-conditional quantum advantage taxonomy. We provide empirical evidence that quantum enhancement benefits specific signal classes—chaotic dynamics, noisy oscillations, and smooth trends—while classical Transformers remain superior for clean periodic signals and sharp discontinuities. This taxonomy offers practitioners a principled guide for when to deploy hybrid quantum-classical models.
  • Trainability and noise resilience of the parsimonious design. Gradient-variance analysis shows that adding a second entangling layer collapses the gradient magnitude by ${\sim}30\times$, approaching a barren plateau—so our single-layer choice is not merely sufficient but necessary for trainability. Under depolarizing noise at hardware-realistic levels ($p\approx 10^{-3}$$10^{-2}$), QASA’s performance remains stable or improves, indicating that the design is NISQ-compatible without requiring error correction.
  • Comprehensive empirical validation. QASA achieves the best MSE on 4 of 9 benchmarks (tied with the classical baseline), significantly outperforms QLSTM on the seasonal trend task (p=0.009p{=}0.009), and reduces MAE by 6.0% on the real-world ETTh1 dataset. These results hold across multiple random seeds and are supported by statistical significance tests, effect size analysis, and non-parametric Friedman tests.
Figure 1: Overview of the QASA architecture. The model consists of N1N{-}1 standard Transformer encoder layers followed by one Quantum Encoder Layer. The Quantum Encoder Layer replaces the value projection in self-attention with a parameterized quantum circuit (PQC) operating on 8 qubits with 4 variational layers, followed by a classical feedforward network with residual connections.
@@ 2.1 Deep Learning on Time-Series Data
Deep learning has significantly advanced the modeling and understanding of time-series data across various domains. Recurrent Neural Networks (RNNs) [33], particularly Long Short-Term Memory (LSTM) networks [23] and Gated Recurrent Units (GRUs) [18], have been widely adopted due to their ability to capture long-term temporal dependencies. However, their sequential nature limits parallelization and increases training time. To address this, temporal convolutional networks (TCNs) [27] and Transformer-based architectures [39] have demonstrated superior performance by enabling parallel processing and better long-range dependency modeling.
In industrial and scientific applications, Transformer variants such as the Temporal Fusion Transformer (TFT) [28] and Informer [40] have been applied to multi-horizon prediction tasks with irregular time intervals and exogenous variables. Similar time-series learning techniques have been applied successfully in molecular dynamics and open quantum systems [37, 38, 7].
Despite these advancements, challenges remain in modeling noisy, sparse, or irregularly-sampled sequences, motivating the development of hybrid models that integrate domain-specific priors with general-purpose deep learning architectures.
@@ 2.2 Quantum Deep Learning on Time-Series Data
While classical deep learning has achieved remarkable success in time-series analysis, it often requires extensive computational resources and large-scale data to generalize effectively. With the advent of quantum computing, Quantum Deep Learning (QDL) [3] has emerged as a promising paradigm for modeling complex temporal patterns, offering theoretical advantages in expressivity and optimization through quantum entanglement and superposition.
In the context of time-series data, several quantum-inspired architectures have been proposed. Variational Quantum Circuits (VQCs) [20] and Quantum Recurrent Neural Networks (QRNNs) [35] have been introduced to capture temporal dependencies using parameterized quantum gates and recurrent structures. A particularly notable development is the Quantum Long Short-Term Memory (QLSTM) network [15], which integrates quantum circuits into the gating mechanisms of classical LSTMs. Recent comprehensive benchmarking studies [19, 9] have compared VQC-based regression against classical GRU and LSTM models, finding that while classical models often lead in raw accuracy, VQCs exhibit higher robustness to noise and better generalization on small-scale datasets. Quantum-enhanced channel mixing has also been explored in state-space models for temporal forecasting [11], while quantum temporal convolutional networks have been applied to cross-sectional equity return prediction [14].
Recent work has further expanded quantum approaches for temporal data. Tang et al. [36] proposed QnnFormer, a multi-attention mechanism leveraging quantum circuits for long-term forecasting, achieving significant error reductions on benchmark datasets. Chakraborty and Heintz [6] integrated hybrid quantum-classical attention into patch-based transformers for multivariate forecasting. Chittoor et al. [17] developed QuLTSF for long-term forecasting with quantum neural networks, demonstrating advantages in convergence stability for non-stationary data. Laskar and Goel [26] investigated the impact of different entanglement topologies in shallow circuits for time-series prediction on NISQ hardware. In finance, quantum-enhanced architectures have been investigated for demand prediction and volatility modeling [32].
Beyond time series, quantum machine learning has been applied to sequential data in other domains, including EEG encoding [8, 10] and multimodal contrastive learning [13]. Chen and Kuo [12] demonstrated that Transformer-assisted learning can effectively capture Lindblad dynamics in open quantum systems, establishing a direct connection between Transformer architectures and quantum physics. Despite these advancements, research on Quantum Transformers for time-series data remains limited. Most existing studies focus on Quantum Visual Transformers (QViT) [16] or fully quantum architectures such as Quixer [25] in sequence modeling, leaving a gap in understanding how hybrid quantum-classical self-attention mechanisms can be practically deployed for temporal forecasting. Addressing this gap is the key focus of this work.
@@ 3 Methodology
Figure 2: The parameterized quantum circuit (PQC) used in QASA’s QuantumLayer. The circuit operates on n=8n{=}8 data qubits plus one ancilla qubit, with L=4L{=}4 variational layers. Each layer applies single-qubit rotations ($R_{X}$, $R_{Y}$, $R_{Z}$; see Eq. (19) for definitions) and CNOT entangling gates in a ring topology. Input features are encoded via $R_{X}$ and $R_{Z}$ rotations on each qubit, and the output is obtained by measuring $\langle Z_{i}\rangle$ on all data qubits.
Table 1: Quantum Adaptive Self-Attention (QASA) Transformer Architecture. Each row lists a layer with its operation and input/output tensor shapes. $L$ denotes the input sequence length (e.g., L=50L{=}50), $d$ is the hidden feature dimension (e.g., d=256d{=}256), and $n$ is the number of qubits in the quantum circuit (e.g., n=8n{=}8). ’Attn’ is an abbreviation for ’attention’.
@@
Transformer Layer $\times(N{-}1)$ Multihead Attn + FFN $(L,d)$ $(L,d)$
@@ 3.3 Transformer Encoder Layers
The encoder consists of $N$ layers, where the first N1N-1 layers are standard Transformer encoder layers defined as:
@@
$\text{MultiHeadSelfAttention}(X)=\text{Concat}(head_{1},\dots,head_{H})W^{O},$ (6)
@@ 3.4 Quantum Encoder Layer
$\text{where }head_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}),$ (15)
@@
The vector $h_{q}$ is then passed as input to a parameterized quantum circuit (PQC), which is defined over $n+1$ qubits and $L_{q}$ layers of unitary operations. The PQC applies data re-uploading using $R_{X}$ and $R_{Z}$ gates, and introduces entanglement using CNOT and $R_{Y}$ gates. Throughout this paper, the single-qubit rotation gates are defined as
$R_{X}(\theta)=e^{-i\theta X/2},\quad R_{Y}(\theta)=e^{-i\theta Y/2},\quad R_{Z}(\theta)=e^{-i\theta Z/2},$ (19)
where $X,Y,Z$ are the Pauli matrices. The final output is computed as the expectation values of Pauli-$Z$ operators on the first $n$ qubits:
@@
The resulting vector $h_{q}$ is fed into a parameterized quantum circuit $\text{QC}(\cdot)$ with $L_{q}$ layers and $n+1$ qubits, implemented using a data reuploading strategy. Each layer of the PQC applies a combination of single-qubit rotations ($R_{X}$, $R_{Y}$, $R_{Z}$) and entangling CNOT gates, allowing the circuit to perform complex non-classical transformations:
@@
In our design, we adopt a hybrid encoder composed of (N1)(N{-}1) standard Transformer encoder layers followed by a single quantum-enhanced encoder layer. This layered configuration offers a practical and effective trade-off between scalability and expressiveness. Specifically, the classical Transformer blocks serve as powerful hierarchical feature extractors with stable training dynamics, while the quantum encoder layer provides a complementary inductive bias via entangled quantum operations and high-dimensional projection.
@@
Training via the parameter-shift rule.
Gradients of the PQC output with respect to its variational parameters are computed using the parameter-shift rule [29]. For any Pauli rotation gate $R_{\sigma}(\theta)=e^{-i\theta\sigma/2}$ with $\sigma\in\{X,Y,Z\}$, and any observable $\hat{O}$, the partial derivative of the expectation value $\langle\hat{O}\rangle_{\theta}=\langle\psi(\theta)|\hat{O}|\psi(\theta)\rangle$ satisfies
$\frac{\partial\langle\hat{O}\rangle_{\theta}}{\partial\theta}=\frac{1}{2}\left[\langle\hat{O}\rangle_{\theta+\pi/2}-\langle\hat{O}\rangle_{\theta-\pi/2}\right].$ (26)
This identity yields exact (noise-free in the ideal case) gradients through two additional circuit evaluations per parameter, and is compatible with PyTorch’s autograd via PennyLane, enabling end-to-end training of the full hybrid model.
QASA forward pass.
Algorithm 1 summarises the forward pass of QASA for a single input sequence, making the hybrid classical–quantum information flow explicit.
Algorithm 1 QASA forward pass for a single input sequence.
1:Input sequence $x\in\mathbb{R}^{L\times 1}$; hidden dim $d$; encoder depth NN; qubit count $n$; PQC depth $L_{q}$
2:Scalar prediction $\hat{y}\in\mathbb{R}$
[…]
4:for $i=1$ to N1N-1 do $\triangleright$ N1N{-}1 classical Transformer layers
5:  $z_{i}\leftarrow\mathrm{LayerNorm}\!\left(h_{i-1}+\mathrm{MultiHeadSelfAttn}(h_{i-1})\right)$
6:  $h_{i}\leftarrow\mathrm{LayerNorm}\!\left(z_{i}+\mathrm{FFN}(z_{i})\right)$
7:end for
[…]
10:for each token $h^{\prime}_{j}$ in the sequence do
11:  $h_{q}\leftarrow\tanh(W_{q}h^{\prime}_{j})+t$ $\triangleright$ classical-to-quantum projection
12:  $o_{j}\leftarrow\mathrm{PQC}_{\theta}(h_{q})$ where $(o_{j})_{k}=\langle Z_{k}\rangle$ for $k=1,\dots,n$ $\triangleright$ Eq. 19
13:  $z_{j}\leftarrow h^{\prime}_{j}+W_{o}o_{j}$ $\triangleright$ residual quantum-to-classical projection
14:end for
[…]
17:return $\hat{y}$
@@
(d) $QASA_{classical}$ - Epoch 1
(e) $QASA_{classical}$ - Epoch 15
(f) $QASA_{classical}$ - Epoch 30
(g) QASAQASA - Epoch 1
(h) QASAQASA - Epoch 15
(i) QASAQASA - Epoch 30
@@
    Data Encoding via Single-Qubit Rotations: The input features are encoded into quantum states using $R_{X}$ and $R_{Z}$ gates per qubit:
[…]
@@
    Parameterized Rotations with Data Reuploading: Across $L_{q}$ layers, the circuit includes learnable $R_{Y}$ and $R_{Z}$ rotations:
    $R_{Y}(\theta_{l,i})\ R_{Z}(\theta_{l,i})\quad\forall i,\forall l\in\{1,\dots,L_{q}\},$ (28)
@@
    $\text{CNOT}(i\rightarrow(i+1)\bmod n),\quad\text{for }i=0,\dots,n{-}1.$ (29)
@@
    $\text{CNOT}(n{-}1\rightarrow n),\quad R_{Y}(\theta_{l,n})\text{ on wire }n.$ (30)
This gate configuration balances expressiveness and hardware feasibility. The use of simple universal gates ($R_{X},R_{Y},R_{Z}$, CNOT) ensures compatibility with most current quantum hardware, while circular entanglement introduces full connectivity with only $n$ CNOTs per layer, avoiding unnecessary depth. The auxiliary $(n+1)$-th qubit also serves as a tunable global channel, useful for capturing long-range dependencies in the encoded features.
3.5 Circuit Expressibility and Entanglement Analysis
To characterize the representational capacity of the QASA quantum circuit beyond its architectural description, we evaluate three established metrics from the quantum circuit benchmarking literature: entanglement entropy, the Meyer–Wallach entangling capability measure, and expressibility.
Entanglement Entropy.
We compute the bipartite von Neumann entropy $S(\rho_{A})=-\mathrm{tr}(\rho_{A}\log_{2}\rho_{A})$ across a balanced bipartition of the data qubits (qubits 0–3 vs. 4–7 plus auxiliary), averaged over 200 random parameter samples. The QASA circuit achieves a mean entanglement entropy of $2.26\pm 0.12$ bits out of a maximum of 4.0 bits (normalized: 0.57), indicating substantial but not maximal entanglement. This moderate level is desirable: it ensures non-trivial quantum correlations across qubit subsets without collapsing into a maximally entangled state that would limit trainability [30].
Meyer–Wallach Entangling Capability.
The Meyer–Wallach measure [31] quantifies the global entanglement of a pure state $|\psi\rangle$ on $N$ qubits as
$Q(|\psi\rangle)=2-\frac{2}{N}\sum_{k=1}^{N}\mathrm{tr}\!\left(\rho_{k}^{2}\right),$ (31)
where $\rho_{k}=\mathrm{tr}_{\neg k}\!\left(|\psi\rangle\langle\psi|\right)$ is the reduced density matrix of qubit $k$. By construction, $Q\in[0,1]$: $Q=0$ for product states and $Q=1$ when every single-qubit reduced density matrix is maximally mixed. Averaged over 1,000 random parameterizations, QASA achieves $Q=0.981$, substantially higher than QLSTM ($Q=0.710$) and QnnFormer ($Q=0.883$). This confirms that the ring-topology CNOT structure with auxiliary qubit coupling generates near-maximally entangled states, despite using fewer two-qubit gates than QLSTM (27 vs. 56 CNOTs).
Expressibility.
Following Sim et al. [34], we quantify the expressibility of a PQC $U(\theta)$ as the Kullback–Leibler divergence between the distribution of pairwise state fidelities generated by random circuit parameters and the Haar-random fidelity distribution:
$D_{\mathrm{KL}}=D_{\mathrm{KL}}\!\left(\,P_{\mathrm{PQC}}(F)\,\|\,P_{\mathrm{Haar}}(F)\,\right),\qquad F=\left|\langle\psi(\theta_{1})\,|\,\psi(\theta_{2})\rangle\right|^{2},$ (32)
where the Haar reference on $N$ qubits is $P_{\mathrm{Haar}}(F)=(d-1)(1-F)^{d-2}$ with $d=2^{N}$. We sample 5,000 random parameter pairs $(\theta_{1},\theta_{2})$ and estimate $P_{\mathrm{PQC}}$ via a 75-bin histogram. QASA yields $D_{\mathrm{KL}}=0.029$, comparable to QnnFormer ($D_{\mathrm{KL}}=0.026$) and substantially better than QLSTM ($D_{\mathrm{KL}}=0.125$). Lower $D_{\mathrm{KL}}$ indicates wider coverage of the Hilbert space, a necessary condition for effective variational learning.
Table 2: Comparative quantum circuit analysis across all three quantum models (5,000 parameter samples for expressibility, 1,000 for entangling capability). Lower expressibility KL divergence indicates wider Hilbert space coverage; higher Meyer–Wallach $Q$ indicates stronger entanglement generation.
Circuit Qubits Q-Params CNOT Gates Total Gates Expr. ($D_{\mathrm{KL}}\downarrow$) Ent. Cap. ($Q\uparrow$)
QASA 9 36 27 110 0.029 0.981
QLSTM 8 32 56 104 0.125 0.710
QnnFormer 8 24 21 94 0.026 0.883
Table 2 reveals a striking efficiency gap: QASA achieves the highest entangling capability ($Q=0.981$) with only 27 CNOT gates, while QLSTM requires more than twice as many CNOTs (56) yet produces the weakest entanglement ($Q=0.710$). This demonstrates that QASA’s ring topology—where each qubit is entangled with its neighbor in a circular arrangement—is fundamentally more efficient than QLSTM’s linear chain topology. QnnFormer achieves competitive expressibility with the fewest total gates (94), but its entangling power remains below QASA’s. These results provide a circuit-level explanation for QASA’s strong empirical performance: near-maximal entanglement enables the quantum layer to capture complex correlations, while high expressibility ensures the circuit can represent diverse functions during training.